OpenStudy (khally92):

Solve the system of the linear equation and interpret your solution geometrically 2x+y+2z-4=0 x-y-z-2=0 x+2y-6z-12=o. I have solved so many, am finding this one tricky, we use the elimination method to eliminate one of the variables..

OpenStudy (cwrw238):

no. 3 - no. 2 gives -3y + 5z + 10 = 0 no. 1 - 2* no 2 gives 3y + 4z = 0 adding the above 9z = -10 z = -10/9 - can you continue with this (i'm tired!!!)

OpenStudy (khally92):

let me solve ill get back to you very soon.

OpenStudy (cwrw238):

ok - the -10/9 is correct - i checked it out on my calculator

OpenStudy (khally92):

check no3-no 2 u made a mistake with the sign

OpenStudy (cwrw238):

lol - sorry - although i put 3 - 2 - did 2 - 3 so it is correct

OpenStudy (khally92):

which should i do?

OpenStudy (cwrw238):

2 - 3 is ok

OpenStudy (khally92):

2-3 after which i will do 1-2 rty?

OpenStudy (cwrw238):

now you you could plug z = -10/9 into equations 1 and 2 and solve for x and y

OpenStudy (radar):

cwrw238 has correct value for z. This problem does not have pretty answers. But you are on track with cwrw238.

OpenStudy (khally92):

did you say 1-2? for the second one

OpenStudy (khally92):

thanks radar?

OpenStudy (khally92):

am confused????

OpenStudy (khally92):

i did 2-3 i got -3y+5z+10=0 what should i do next???

OpenStudy (cwrw238):

no i multiplied no 2 by 2 because no 1 has 2x in it so no 2 becomes 2x - 2y - 2z - 4 = 0 im afraid i didn't show this step then no 1 - this equation gave 3y + 4z = 0

OpenStudy (cwrw238):

-3y+5z+10=0 is correct now add 3y + 4z = 0 (derived above) to get an equation in z

OpenStudy (khally92):

good i now have the two equations so i should add them up right?

OpenStudy (cwrw238):

yes add them

OpenStudy (khally92):

yeah z= -10/9

OpenStudy (cwrw238):

right now you can plug z into, i suggest, no 1 and no 2 equations to give 2 equations in x and y

OpenStudy (khally92):

for the first one 2x+y-56/9=0

OpenStudy (khally92):

and for the second one x-y-8/9=0

OpenStudy (cwrw238):

yes

OpenStudy (khally92):

yeah so whats next?

OpenStudy (cwrw238):

well solve these 2 equations by elimination adding them will eliminate y ( as they contain y and -y)